NatalieBraudt:

Fourteen percent of the town's population is over the age of 65. If there are 320 residents over the age of 65, approximately what is the town's population? 45 5 2,286 2,564

Hoodmemes:

Honestly first I think u need to do/know that 14 percent of the towns population that can be rewritten as 14/100 times the Town's population = 320. Then solve 14/100 Population = 320,Multiply both sides by the reciprocal of 14/100. Then Multiply by 100/14. So Town population = 320 times 100/14 = 32000/14=2285.71 or approximately= (blank) people. (do u kinda understand it?)

Hoodmemes:

Close this question then post your new question.